EVA Report – January 23rd

Crew 220 EVA #5 Report 23-01-2020

EVA # 5

Author: Johannes Svensoy, MD

Purpose of EVA: Recognition for sites usable for Medical Trauma training.

Start time: 0900

End time: 1200

Narrative: Pre-EVA: All EVA Crew members physically fit, hydrated, and mentally prepared. Crew was briefed about mission for finding two sites which will be good for medical trauma training. Preferably around the MAU station. Terrain muddy. Extra caution was taken throughout the mission.

Airlock working perfectly, and after a successful depressurization the Crew drove over to MAU station with the Spirit and Curiosity rovers. At the MAU station, the MDRS Crew met up with the MAU Crew. After a quick brief and visual recognition of the surrounding area, the whole group ventured south west. The first site was marked at 12s518124,4250680. From there another site were located west of the MAU station at 12s518287,4250596. Several places around the second site were identified as potential good places for training which is of major importance for the Crew for future possible medical traumas here on Mars.

All astronauts returned safely to MDRS with no issues. They all went through post-EVA medical checkup with no issues found.

Destination: Recognition of area South of MDRS

Coordinates (use UTM NAD27 CONUS):

Participants: HabCom: Lee. EVA Crew: Shawna, Johannes, Susan.

Road(s) and routes per MDRS Map: Entrance road out to Cow Dung Road. Area around MAU station.

Mode of travel: Rovers Spirit and Curiosity from MDRS to Cow Dung Road. Parked at MAU station. Trek around the MAU station area.

Support Requested: No support requested during EVA.

Spirit rover used:

Beginning Hours: 127.7

Ending Hours: 127.8

Beginning charge: 100%

Ending charge: 88%

Currently charging: YES

Curiosity rover used:

Beginning Hours: 133.7

Ending Hours: 133.7

Beginning charge: 100%

Ending charge: 87%

Currently charging: YES

GreenHab Report – January 23rd

Crew 220 GreenHab Report 23-JAN-20

Crew GreenHab Officer: Marufa Bhuiyan

Average Temperature (Floor Unit): 14°C

Floor temperature at different times:
09:30: 20°C
15:08: 8°C

Environmental Heating:
GreenHab door Shade cloth (40%) (30%) on
Door opened: 12:40
Door closed: 15:08

Time(s) of watering crops:

Daily water usage for crops: 11 Gal
Morning: 5 Gal
Evening: 6 Gal

Daily water usage for research and/or other purposes: N/A

Water in the Blue Tank: 211 Gal

Harvest: N/A

Changes in Crops/Crops picked for cooking purpose: Chives, Basil.

Other comments: N/A

Operations Report – January 23rd


 Crew 220 Operations Report
SOL: 4

Submitting Report: Lee Roberts, Crew Engineer

Non-nominal systems: Primary generator and water heater are inoperative

Generator: Primary generator is offline, power system operating on rental generator

Hours Run: 7:30

Start Time: Ran through the night

End Time: 0730

Solar SOC: 82%

Diesel Reading: 49%

Station Propane Reading: 90%

Ethanol Free Gasoline: N/A

Water in Loft Tank: 50 Gallons

Water Meter: 148107.4

Water in Static Tank: 352 Gallons

Static to Loft Tank: 43 Gallons

Water in GreenHab Tank: 211 Gallons

Water in Science Dome: N/A

Toilet Tank Emptied: Yes

Spirit Rover Used: Yes

Ending Hours: 127.9

EVA 1 Beginning Charge: 100%

EVA 1 Ending Charge: 88%

EVA 2 Beginning Charge: 100%

EVA 2 Ending Charge: 74%

Curiosity Rover Used: Yes

Ending Hours: 133.8

EVA 1 Beginning Charge: 100%

EVA 1 Ending Charge: 87%

EVA 2 Beginning Charge: 100%

EVA 2 Ending Charge: 72%

Notes on Rovers: None

ATVs Used: None

HabCar Used, Why, Where? Negative

Crew Car Used, Why, Where? Negative
General Notes and Comments: None

Summary of Internet: Internet is operating nominally

Summary of Suits and Radios: Communications were intermittent on two of the radios. Spent the afternoon troubleshooting.

Summary of Hab Operations: With noted exceptions, hab systems are operating nominally. The ventilation system air filter was replaced this morning.

Summary of GreenHab Operations: The automatic timer on the lights was out of sync due to the power outage of the last two days. I am resetting the time to correspond to the local clock time. Please refer to the GreenHab Report for further details.

Summary of RAM Operations: An air filter was taken from the RAM supply closet to replace the ventilation system air filter.

Summary of Observatory Issues: None

Summary of Health and Safety Issues: Please refer to the Health and Safety Report

Questions, concerns, and requests to Mission Support: We have no specific concerns or requests at this time

Sol Summary – January 23rd

MDRS Crew 220 Sol 4 Summary 23Jan2020

Summary Title: Target acquired?
Author’s name: Shawna Pandya, Commander
Mission Status: Nominal
Sol Activity Summary: Today was a busy day as both crews ran joint EVA exercises in the morning and afternoon with the goal of scouting injury sites for future medical simulations on EVA, followed by a tour of the MAU station. The morning EVA went smoothly, with 2 injury sites (IS) easily found and demarcated.. The afternoon EVA teams, based on the instructions of the morning EVA team found both IS, but then had to end their EVA early due to several comms failure. All in all, the morning and afternoon EVAs achieved the desired objectives, setting the stage for future medical EVAs. Within the Hab, the team continued to bring the surveys, tech and science back online that had previously been disrupted with the loss of Internet.Look Ahead Plan: Tomorrow’s EVAs will consist of medical drills at the injury sites identified today and briefing on medical triage, assessment and communication protocols in the field.
Anomalies in work: Hot water still out.
Weather: -5C low / 2C max, 0C average, 20% chance of precipitation tomorrow
Crew Physical Status: Excellent
EVA: 1 AM MDRS EVA and 1 PM MDRS EVA performed today as above
Reports to be filed:
Sol Summary
Journalist’s Report
Ops/Engineering Report
GreenHab Report
Photo Report
EVA Request: To be submitted
Support Requested: Please remove trash from airlock. Requesting saran wrap, aluminum foil, baby wipes, 1 LED lightbult please.

Copyright © The Mars Society. All rights reserved. | Main Site